# A posteriori error estimation for an interior penalty virtual element method of Kirchhoff plates

### v2

@@ 5.2   Lower bound
By means of the technique of bubble functions frequently used in the a posteriori error analysis [21, 24], we can also derive the following lower bound of the a posteriori error estimator.
@@
$h^{4}_{K}\|f_{h}-\Delta^{2}\Pi_{h}^{\Delta}w_{h}\|_{0,K}^{2}$ $\lesssim|u-\Pi_{h}^{\Delta}w_{h}|_{2,K}^{2}+\eta_{5,K}^{2},$ (5.6)
$h_{e}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\Delta}w_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\Delta}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2},$ (5.7)
$h_{e}^{3}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\Delta}w_{h}}{\partial\bm{n}_{e}}+\frac{\partial^{3}\Pi_{h}^{\Delta}w_{h}}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\Delta}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2}.$ (5.8)
@@
We divide the proof into three steps. For brevity, we omit the subscript $e$ in $\bm{t}_{e}$ and $\bm{n}_{e}$. For the sake of brevity, we denote $v=\Pi_{h}^{\Delta}w_{h}$ in what follows.
@@
$\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}^{2}_{0,e}\lesssim\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\partial_{\bm{n}}\phi\Big{)}_{e}$ (5.9)
@@
$\Big{\|}\Big{[}\frac{\partial\Delta v}{\partial\bm{n}}+\frac{\partial^{3}v}{\partial\bm{n}\partial\bm{t}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim\sum\limits_{\tau=\tau^{+},\tau^{-}}\Big{(}(\Delta^{2}v-f,\,\phi)_{\tau}+a^{\tau}(u-v,\,\phi)\Big{)}+\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\partial_{\bm{n}}\phi\Big{)}_{0,e}$
@@
$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}+J_{1}(u-u_{h},u-u_{h})$
@@
By the norm equivalence (3.17),
$|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}$ $\lesssim|u-u_{h}|_{2,K}^{2}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}$
$\lesssim|u-u_{h}|_{2,K}^{2}+S^{K}(u_{h}-\Pi_{h}^{\Delta}u_{h},\,u_{h}-\Pi_{h}^{\Delta}u_{h})$
@@
It follows from the norm equivalence in (3.17) and the triangle inequality that
$\eta_{4,K}$ $=h_{K}^{-1}\|\bm{\chi}(u_{h}-\Pi_{h}^{\Delta}u_{h})\|_{\ell^{2}}\eqsim|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}\lesssim|u-u_{h}|_{2,K}+|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}.$
Here, by the triangle inequality, (3.8) and (2.4), we derive
$|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}\leq|\Pi_{h}^{\Delta}u-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|u-\Pi_{h}^{\Delta}u|_{2,K}\lesssim|u-u_{h}|_{2,K}+h^{k-1}.$
Inserting (5.20) and (5.2) into (5.2), one has
$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-u_{h}|_{2,K}^{2}+J_{1}(u-u_{h},u-u_{h})+\mathcal{O}(h^{2k-2}).$ (5.22)
@@ 5.3   Comparison with the analysis of Morley-type virtual elements
Our analysis of the upper bound in Theorem 5.1 can be extended to the Morley-type VEM. Utilizing Lemma 5.1, we can avoid the tedious integration by parts for the term $I_{2}$ as follows:
@@
Another difference in the error estimators from [13] is the substitution of the local estimators $\eta_{0,e}(u_{h})$ and $\eta_{1,e}(u_{h})$ with $\eta_{0,e}(\Pi_{h}^{\Delta}u_{h})$ and $\eta_{1,e}(\Pi_{h}^{\Delta}u_{h})$. For the Morley-type VEM in [13], unlike the triangle inequality
@@
$|u-u_{h}|_{2,K}\leq|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-u_{h}|_{2,K}.$ (5.24)
The second term corresponds to the error estimator $\eta_{4,K}$. For the first term, since $\Pi_{h}^{\Delta}u_{h}|_{K}$ is a polynomial on $K$, they treat $\Pi_{h}^{\Delta}u_{h}$ as a finite element function on the virtual triangulation $\widehat{\mathcal{T}}$, defining $\widehat{E}_{h}:\mathbb{P}_{2}(\widehat{\mathcal{T}})\to H_{0}^{2}(\Omega)$ as the enriching operator for the FEM. They then decompose the first term as
$|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}\leq|u-\widehat{E}_{h}\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-\widehat{E}_{h}\Pi_{h}^{\Delta}u_{h}|_{2,K}.$ (5.25)
Let $\widehat{E}_{h}\Pi_{h}^{\Delta}:=\widetilde{E}_{h}$. The decomposition given by (5.24) and (5.25) can be written as
@@
$\leq|u-\widetilde{E}_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-\widetilde{E}_{h}u_{h}|_{2,K}.$
The first term can be estimated as for the first term in (5.23), namely the first term in (5.2) for Theorem 5.1. By using the standard result for $\widehat{E}$, the third term is bounded by
$|\Pi_{h}^{\Delta}u_{h}-\widetilde{E}_{h}u_{h}|_{2,h}^{2}=|\Pi_{h}^{\Delta}u_{h}-\widehat{E}_{h}\Pi_{h}^{\Delta}u_{h}|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\Delta}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
For any $v\in V_{h}$, it is evident that $E_{h}\phi$ can be defined for $\phi=\Pi_{h}^{\Delta}v$. By the definition of $E_{h}$, we have $E_{h}v=E_{h}\Pi_{h}^{\Delta}v$, which along with Lemma 4.1 implies
$|\Pi_{h}^{\Delta}v-E_{h}v|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\Delta}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
@@
$\leq|u-E_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\Delta}u_{h}|_{2,K}+|\Pi_{h}^{\Delta}u_{h}-E_{h}u_{h}|_{2,K}.$
This approach allows us to derive the error estimators in [13] for the Morley VEMs of arbitrary order, not just the lowest order. However, for interior penalty methods, this may not be appropriate, as including the elliptic projector $\Pi_{h}^{\Delta}$ in $J_{1}$ (see (3.11)) is not feasible because $\|\cdot\|_{h}$ may not be a norm on $V_{h}$ in this context and $J_{1}(u-u_{h},u-u_{h})$ will introduce the a-priori information.
@@ 6 Numerical experiments
In this section, we present several benchmark tests for which the exact solution is known or at least its behavior is known in a qualitative sense. These tests illustrate the robustness of the residual-type a posteriori error estimator and demonstrate the efficiency of the adaptive VEM. Unless otherwise specified, we only consider the lowest-order element $k=2$. Since the VEM solution $u_{h}$ is not explicitly known inside the polygonal elements, we evaluate the projection error, which is quantified by
[…]
The first term in (3.9) will be disregarded when focusing solely on the $H^{2}$-semi-norm.
@@
All examples are implemented in MATLAB R2019b. Our code is available from GitHub (https://github.com/Terenceyuyue/mVEM) as part of the mVEM package which contains efficient and easy-following codes for various VEMs published in the literature. The subroutine IPVEM_Zhao.m is used to compute the numerical solutions. The test script main_IPVEM_Zhao_avem.m verifies the convergence rates. The error estimator is computed by the subroutine IPVEM_Zhao_indicator.m.
@@ 6.2   Solution with large gradients
Fig. 4: The initial and the final adapted meshes. (a) The initial mesh; (b) After 20 refinement steps; (c) After 30 refinement steps; (d) The zoomed mesh in (c)
@@
Fig. 6: The exact and numerical solutions
For this example, we choose the parameter $\theta=0.6$. The initial mesh and the final adapted meshes after 20 and 30 refinement steps are presented in Fig. 4 (a-c), respectively. The detail of the last mesh is shown in Fig. 4 (d). Clearly, no small edges are observed. We plot the adaptive convergence order in Fig. 6, which confirms the theoretical prediction as in the last example. We also display the numerical and exact solutions in Fig. 5, from which we see the adaptive strategy correctly refines the mesh in a neighborhood of the singular point and there is a good level of agreement between the $H^{2}$ error and error estimator.
7 Conclusion
We developed a residual-type a posteriori error estimator for an interior penalty virtual element method applied to a Kirchhoff plate bending problem. A key step involves constructing an enriching operator to establish the reliability and efficiency of the a posteriori error bound. We presented an error estimate for the enriching operator that aligns with the interior penalty finite element method, as detailed in Lemma 4.1. Additionally, we provided an example estimate in Lemma 4.2 with a proof idea applicable to the medius error analysis.

### v4

@@ 5.2   Lower bound
By means of the technique of bubble functions frequently used in the a posteriori error analysis [21, 24], we can also derive the following lower bound of the a posteriori error estimator. The proof is similar to [21], but for the sake of completeness, we still provide a detailed proof.
@@
$h^{4}_{K}\|f_{h}-\Delta^{2}\Pi_{h}^{\nabla}w_{h}\|_{0,K}^{2}$ $\lesssim|u-\Pi_{h}^{\nabla}w_{h}|_{2,K}^{2}+\eta_{5,K}^{2},$ (5.7)
$h_{e}\Big{\|}\Big{[}\frac{\partial^{2}\Pi_{h}^{\nabla}w_{h}}{\partial\bm{n}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\nabla}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2},$ (5.8)
$h_{e}^{3}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}w_{h}}{\partial\bm{n}_{e}}+\frac{\partial^{3}\Pi_{h}^{\nabla}w_{h}}{\partial\bm{n}_{e}\partial\bm{t}_{e}^{2}}\Big{]}\Big{\|}_{0,e}^{2}$ $\lesssim|u-\Pi_{h}^{\nabla}w_{h}|_{2,\omega(e)}^{2}+\sum_{K^{{}^{\prime}}\in\omega(e)}\eta_{5,K^{{}^{\prime}}}^{2}.$ (5.9)
@@
We divide the proof into three steps. For brevity, we omit the subscript $e$ in $\bm{t}_{e}$ and $\bm{n}_{e}$. For the sake of brevity, we denote $v=\Pi_{h}^{\nabla}w_{h}$ in what follows.
@@
$\Big{\|}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]}\Big{\|}^{2}_{0,e}\lesssim\Big{(}\Big{[}\frac{\partial^{2}v}{\partial\bm{n}^{2}}\Big{]},\partial_{\bm{n}}\phi\Big{)}_{e}.$ (5.10)
@@
[…]
@@
$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}^{2}+J_{1}(u-u_{h},u-u_{h})$
@@
By the inverse inequality (3.6) and the norm equivalence (3.20),
$|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}^{2}$ $\lesssim|u-u_{h}|_{2,K}^{2}+|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}^{2}$
$\lesssim|u-u_{h}|_{2,K}^{2}+h_{K}^{-2}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{1,K}^{2}$
@@
It follows from the norm equivalence in (3.20), the inverse inequality (3.6) and the triangle inequality that
$\eta_{4,K}$ $=h_{K}^{-1}\|\bm{\chi}(u_{h}-\Pi_{h}^{\nabla}u_{h})\|_{\ell^{2}}\eqsim h_{K}^{-1}|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{1,K}\lesssim|u-u_{h}|_{2,K}+|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}.$
Here, by the triangle inequality, Lemma 3.3 and (2.4), we derive
$|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}\leq|\Pi_{h}^{\nabla}u-\Pi_{h}^{\nabla}u_{h}|_{2,K}+|u-\Pi_{h}^{\nabla}u|_{2,K}\lesssim|u-u_{h}|_{2,K}+h^{k-1}.$
From the continuity of $u$, the trace inequality and (2.4), we get
$J_{1}(u-u_{h},u-u_{h})$ $\lesssim\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial(\Pi_{h}^{\nabla}u-u)}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}+\sum_{e\in\mathcal{E}_{h}}\frac{1}{|e|}\Big{\|}\Big{[}\frac{\partial\Pi_{h}^{\nabla}u_{h}}{\partial\bm{n}_{e}}\Big{]}\Big{\|}_{0,e}^{2}$
$\lesssim\sum_{K\in\mathcal{T}_{h}}|\Pi_{h}^{\nabla}u-u|_{2,K}^{2}+h_{K}^{-2}|\Pi_{h}^{\nabla}u-u|_{1,K}^{2}+J_{1}(u_{h},u_{h})$
$\lesssim h^{2k-2}+J_{1}(u_{h},u_{h}).$ (5.23)
Inserting (5.21),(5.2) and (5.2) into (5.2), one has
$\eta^{2}(u_{h})\lesssim$ $\sum_{K\in\mathcal{T}_{h}}|u-u_{h}|_{2,K}^{2}+J_{1}(u_{h},u_{h})+\mathcal{O}(h^{2k-2}).$ (5.24)
@@
Remark 5.1.
In the lower bound estimate, the error estimates for $u-\Pi_{h}^{\nabla}u$ are not the a priori estimates for the discrete problem.
@@ 5.3   Comparison with the analysis of Morley-type virtual elements
Our analysis of the upper bound in Theorem 5.1 can be extended to the Morley-type VEM. Using Lemma 5.1, we can avoid tedious integration by parts for the term $I_{2}$ as follows:
@@
For the Morley-type VEM in [13], unlike the triangle inequality
@@
$|u-u_{h}|_{2,K}\leq|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}+|\Pi_{h}^{\nabla}u_{h}-u_{h}|_{2,K}.$ (5.26)
The second term corresponds to the error estimator $\eta_{4,K}$. For the first term, since $\Pi_{h}^{\nabla}u_{h}|_{K}$ is a polynomial on $K$, they treat $\Pi_{h}^{\nabla}u_{h}$ as a finite element function on the virtual triangulation $\widehat{\mathcal{T}}$, defining $\widehat{E}_{h}:\mathbb{P}_{2}(\widehat{\mathcal{T}})\to H_{0}^{2}(\Omega)$ as the enriching operator for the FEM. They then decompose the first term as
$|u-\Pi_{h}^{\nabla}u_{h}|_{2,K}\leq|u-\widehat{E}_{h}\Pi_{h}^{\nabla}u_{h}|_{2,K}+|\Pi_{h}^{\nabla}u_{h}-\widehat{E}_{h}\Pi_{h}^{\nabla}u_{h}|_{2,K}.$ (5.27)
Let $\widetilde{E}_{h}=\widehat{E}_{h}\Pi_{h}^{\nabla}$. The decomposition given by (5.26) and (5.27) can be written as
@@
$\leq|u-\widetilde{E}_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}+|\Pi_{h}^{\nabla}u_{h}-\widetilde{E}_{h}u_{h}|_{2,K}.$
The first term can be estimated as for the first term in (5.25), namely the first term in (5.2) for Theorem 5.1. By using the standard result for $\widehat{E}_{h}$, the third term is bounded by
$|\Pi_{h}^{\nabla}u_{h}-\widetilde{E}_{h}u_{h}|_{2,h}^{2}=|\Pi_{h}^{\nabla}u_{h}-\widehat{E}_{h}\Pi_{h}^{\nabla}u_{h}|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\nabla}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
For any $v\in V_{h}$, it is evident that $E_{h}\phi$ can be defined for $\phi=\Pi_{h}^{\nabla}v$. By the definition of $E_{h}$, we have $E_{h}v=E_{h}\Pi_{h}^{\nabla}v$, which along with Lemma 4.1 implies
$|\Pi_{h}^{\nabla}v-E_{h}v|_{2,h}^{2}\lesssim\sum_{e\in\mathcal{E}_{h}}\Big{(}\frac{1}{|e|^{3}}\|[\Pi_{h}^{\nabla}v]\|_{0,e}^{2}+\frac{1}{|e|}\Big{\|}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{\|}_{0,e}^{2}\Big{)}.$
@@
$\leq|u-E_{h}u_{h}|_{2,K}+|u_{h}-\Pi_{h}^{\nabla}u_{h}|_{2,K}+|\Pi_{h}^{\nabla}u_{h}-E_{h}u_{h}|_{2,K}.$
This approach allows us to directly derive the error estimators in [13] for the Morley VEMs of arbitrary order.
@@ 6 Numerical experiments
In this section, we present several benchmark tests for which the exact solution is known. These tests illustrate the robustness of the residual-type a posteriori error estimator and demonstrate the efficiency of the adaptive VEM. Unless otherwise specified, we only consider the lowest-order element $k=2$. We also provide the treatment of inhomogeneous boundary value conditions.
6.1   Treatment of inhomogeneous boundary value conditions
Consider the fourth-order problem with inhomogeneous boundary value conditions
$\left\{\begin{array}[]{ll}\Delta^{2}u=f&{\rm in}~{}~{}\Omega,\\ u=g_{D},\quad\dfrac{\partial u}{\partial\bm{n}}=g_{N}&{\rm on}~{}~{}\partial\Omega,\end{array}\right.$ (6.1)
where $g_{D}\in H^{3/2}(\partial\Omega)$ and $g_{N}\in H^{1/2}(\partial\Omega)$ are prescribed functions. Let $v\in H^{2}(K)$. The integration by parts gives
$(f,v)_{K}=(\Delta^{2}u,v)_{K}=(\nabla^{2}u,\nabla^{2}v)_{K}$
$\hskip 85.35826pt+\int_{\partial K}\Big{(}\frac{\partial\Delta u}{\partial\bm{n}_{K}}v+\frac{\partial^{2}u}{\partial\bm{n}_{K}\partial\bm{t}_{K}}\frac{\partial v}{\partial\bm{t}_{K}}\Big{)}\mathrm{d}s-\int_{\partial K}\frac{\partial^{2}u}{\partial\bm{n}_{K}^{2}}\frac{\partial v}{\partial\bm{n}_{K}}\mathrm{d}s.$
Using the formula in (2.2) and noting that $u\in H^{3}(\Omega)$, we get
yielding
To ensure the symmetry and enforce the $C^{1}$-continuity of the discrete variational problem, we rewrite it as
$(f,v_{h})+\sum_{e\in\mathcal{E}_{h}^{\partial}}\Big{(}-\int_{e}\Big{\{}\frac{\partial^{2}v_{h}}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s+\frac{\lambda_{e}^{3}}{|e|^{3}}\int_{e}[u][v_{h}]\mathrm{d}s+\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial u}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s\Big{)}$
where we have used the fact that $[u]_{e}=0$ and $[\frac{\partial u}{\partial\bm{n}_{e}}]_{e}=0$ for $e\in\mathcal{E}_{h}^{0}$. $\lambda_{e}\geq 1$ is some edge-dependent parameter. Taking into account Remark 3.1, we can describe the IPVEM for (6.1) with inhomogeneous boundary value conditions as follows: Find $u_{h}\in V_{h}^{g}=\{v\in V_{h}:v=I_{h}g_{D}~{}{\rm on}~{}\partial\Omega\}$ such that
$\mathcal{A}_{h}(u_{h},v_{h})=F_{h}(v_{h}),\quad v_{h}\in V_{h},$ (6.2)
where $g_{D}$ has been extended to the whole domain. The right-hand side is
$F_{h}(v_{h})=\langle f_{h},v_{h}\rangle+\sum_{e\in\mathcal{E}_{h}^{\partial}}\int_{e}\Big{(}-g_{N}\frac{\partial^{2}\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}^{2}}+g_{N}\frac{\lambda_{e}}{|e|}\frac{\partial\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}}\Big{)}\mathrm{d}s.$
@@ 6.1   Treatment of inhomogeneous boundary value conditions
All examples are implemented in MATLAB R2019b. Our code is available from GitHub (https://github.com/Terenceyuyue/mVEM) as part of the mVEM package which contains efficient and easy-following codes for various VEMs published in the literature. The subroutine IPVEM.m is used to compute the numerical solutions. The test script main_IPVEM_avem.m verifies the convergence rates. The error estimator is computed by the subroutine IPVEM_indicator.m.
Since the VEM solution $u_{h}$ is not explicitly known inside the polygonal elements, we evaluate the projection error, which is quantified by
[…]
@@ 6.3   Solution with large gradients
Fig. 4: The initial and the final adapted meshes. (a) The initial mesh; (b) After 10 refinement steps; (c) After 20 refinement steps; (d) The zoomed mesh in (c)
@@
Fig. 6: The exact and numerical solutions for Example 6.2
For this example, we choose the parameter $\theta=0.6$. The initial mesh and the final adapted meshes after 10 and 20 refinement steps are presented in Fig. 4 (a-c), respectively. The detail of the last mesh is shown in Fig. 4 (d). Clearly, no small edges are observed. We plot the adaptive convergence order in Fig. 5, which confirms the theoretical prediction as in the last example. We also display the numerical and exact solutions in Fig. 6, from which we see the adaptive strategy correctly refines the mesh in a neighborhood of the singular point and there is a good level of agreement between the $H^{2}$ error and error estimator.
6.4   Solution with low regularity
Example 6.3.
This example examine the performance of the IPVEM under the low regularity condition. The domain is takes as an L-shape square $\Omega=(0,1)^{2}\backslash([1/2,1]\times[0,1/2])$. We select the load term $f$ such that the analytical solution is
$u(x,y)=((x-0.5)^{2}+(y-0.5)^{2})^{5/6}.$
It is evident that the solution $u\in H^{8/3-\epsilon}$ for some $\epsilon>0$ exhibits singular behavior at the point $(1/2,1/2)$. For error evaluation we introduce a small parameter $\delta$ into the exact solution. Specifically, we modify $u$ as $u(x,y)=((x-0.5)^{2}+(y-0.5)^{2}+\delta)^{5/6}$. In this example, we choose the parameter $\theta=0.6$. We plot the adaptive convergence order in Fig. 7, from which we observe that the IPVEM achieves the correct convergence order, even under the low regularity condition of the solution. Additionally, we present the numerical and exact solutions in Fig. 8, which demonstrate that the adaptive strategy effectively refines the mesh in the vicinity of the singular point.
Fig. 8: The exact and numerical solutions for Example 6.3
6.5   Formulation based on $H^{2}$-elliptic projector
In [13, 16], the $H^{1}$-elliptic projector $\Pi_{h}^{\nabla}$ is replaced by the $H^{2}$-elliptic projector $\Pi_{h}^{\Delta}$ in the error estimators. For the IPVEM, we can adopt a similar approach, except for the handling of gradient jumps. In this case, we define
$J_{1}(v,w)=\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big{[}\frac{\partial\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}}\Big{]}\Big{[}\frac{\partial\Pi_{h}^{\Delta}w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
$J_{2}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}v}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s,$
$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{\{}\frac{\partial^{2}\Pi_{h}^{\Delta}w}{\partial\bm{n}_{e}^{2}}\Big{\}}\Big{[}\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big{]}\mathrm{d}s.$
The substitution of the $H^{1}$ projection in the gradient jumps creates difficulty in the process of $I_{2,3}$ in (5.5), as we cannot cancel $J_{2}$. On the other hand, using the $H^{1}$ projection is the preferred choice in the a priori estimate in [18].
We can still prove that $\|w\|_{h}=(|w|_{2,h}^{2}+J_{1}(w,w))^{1/2}$ introduces a norm on $V_{h}$. It is enough to prove that $\|v_{h}\|_{h}=0$ implies $v_{h}=0$ for any given $v_{h}\in V_{h}$. By definition, $\|v_{h}\|_{h}=0$ is equivalent to $|v_{h}|_{2,h}=0$ and $J_{1}(v_{h},v_{h})=0$. Equation $|v_{h}|_{2,h}=0$ shows that $\nabla_{h}v_{h}$ is a piecewise constant on $\mathcal{T}_{h}$. On the other hand, the direct manipulation yields
$\int_{e}\Big{[}\frac{\partial v_{h}}{\partial\bm{n}_{e}}\Big{]}^{2}\mathrm{d}s$ $\leq\int_{e}\Big{[}\frac{\partial\Pi_{h}^{\Delta}v_{h}}{\partial\bm{n}_{e}}\Big{]}^{2}\mathrm{d}s+\int_{e}\Big{[}\frac{\partial(v_{h}-\Pi_{h}^{\Delta}v_{h})}{\partial\bm{n}_{e}}\Big{]}^{2}\mathrm{d}s=\int_{e}\Big{[}\frac{\partial(v_{h}-\Pi_{h}^{\Delta}v_{h})}{\partial\bm{n}_{e}}\Big{]}^{2}\mathrm{d}s,$
$\leq C\sum_{K=K^{-},K^{+}}(|v_{h}-\Pi_{h}^{\Delta}v_{h}|_{1,K}+|v_{h}-\Pi_{h}^{\Delta}v_{h}|_{2,K})\leq C\sum_{K=K^{-},K^{+}}|v_{h}|_{2,K}=0.$
Since $\nabla_{h}v_{h}$ is piecewise constant, we further obtain $[\frac{\partial v_{h}}{\partial\bm{n}_{e}}]=0$ over the edges of $\mathcal{T}_{h}$. That is, the normal derivative of $v_{h}$ is continuous at the interior edges and vanishes at the boundary of $\Omega$. This reduces to the discussion in the proof of Lemma 4.2 in [30], so we omit the remaining argument.
We repeated the test for all the examples and obtained similar results, so we do not include them here. The test script main_IPVEM_avemH2.m verifies the convergence rates. The error estimator is computed by the subroutine IPVEM_indicatorH2.m.